Define: R^3→R^3 by


T(x,y,z)=(x-y+z,x+y,y+z)


Let v1= (1,1,1), v2= (0,1,1), v3= (0,0,1). Find a matrix of T with respect to the basis {v1,v2,v3}. Futher check T is invertible or not.


1
Expert's answer
2021-09-03T02:39:16-0400

Here T("x,y,z)=(x-y+z,x+y,y+z)"

"v_1=(1,1,1)" , "v_2=(0,1,1)" , "v_3=(0,0,1)"

So "T(1,1,1)=(1-1+1,1+1,1+1)=(1,2,2)"

"T(0,1,1)=(0-1+1,0+1,1+1)=(0,1,2)"

"T(0,0,1)=(0-0+1,0+0,0+1)=(1,0,1)"


So Matrix of T with respect to basis {"v_1,v_2,v_3"} is :

"A=\begin{bmatrix}
 1& 2 &2 \\
 0 & 1 & 2\\
 1 &0&1
\end{bmatrix}"


Here |A|="1(1-0)-2(0-2)+2(0-1)=1+4-2=5-2=3" which is not equal to .So matrix is invertible.